TLY operational status:
Telemetry to the station was restored after the station operator visited the station and replaced faulty communications equipment.
Data is latent since day 2020,246-13:17. The station operator reports problems with the satellite modem and transmitter.
The station is back online with an STS-2.5 seismometer and FBA-23 strong motion sensor. The station operators are troubleshooting problems with the original STS-1 seismometers.
The station is down as of day 2015,360-08:52 (Dec. 26) due to power problems. The diesel generator is out of fuel.
The station is back online as of day 2015,294-12:07 (October 21). Power is currently supplied to the station by a diesel generator, however the power system is still unstable with brief power outages.
The station is down as of day 2015,198-11:27 (July 17) due to power problems.
The station is back online as of day 2014,210-04:10.
The station is down as of day 2014,188-16:11 due to a strong storm that passed throught the area. The station operators are investigating this issue.
The East-West channels from the STS-1 showed excessive background noise from days 2013,207-10:47 to 2013,221-08:30 for reasons yet to be determined.
IDA station TLY has been upgraded to a Quanterra Q330HR system. Sensors installed include Steckeisen STS-1 and Teledyne Geotech GS-13 seismometers.
IDA station TLY came back on-line on day 2012,268-07:47. There are intermittent telemetry outages for reasons yet to be determined.
IDA station TLY is down as of day 2012,260-13:05 for reasons yet to be determined.
IDA station TLY is back on-line as of day 2012,235-09:34.
IDA station TLY is down as of day 2012,229-18:57 due to a power outage.
IDA station TLY was down from day 2011,311-03:20 to 2011,314-07:34 due to a power outage.
There may be a failing RT303 board in the DAS. Broadband (32bit) data was not acquired from day 2010,168-17:27 to 2010,171-02:59.
The station has been upgraded to IDA9 data format. The ISP has been upgraded to a SunFire including a CD writer. All the broadband (BB*) channels are seismic as of about day 103-12:22.
Station is up as of day 287-13:57 (Oct.14), except BBZ channels are out for reasons yet to be determined.
Station is down since day 220 (Aug 8) due to problems with DAS. The station operator reported the cause was a thunderstorm August 7.
Telemetry restored. NRTS data coming in as of day 308-09:59 (Nov. 3). Excessive DAS reboots in telemetered and recent tape data indicate the ISP hard disk needs to be replaced.
Station operator reports fixing the station on 2004.03.30. Telemetry is still down due to a broken communication cable.
Station operator reports GPS clock failure on 2003.09.09.
There is frequent spiking on the North-South channels from the STS-1 seismometer for reasons that are not yet understood.
Data acquisition was restored. New EPROMs installed to correct DSP problem.
No data coming from DAS. Failure of DAS XCOM board suspected.
Station operating normally. There is high frequency background noise in the vault.